### v1

@@ Abstract
We consider a problem of inferring contact network from nodal states observed during an epidemiological process. In a black–box Bayesian optimisation framework this problem reduces to a discrete likelihood optimisation over the set of possible networks. The high dimensionality of this set, which grows quadratically with the number of network nodes, makes this optimisation computationally challenging. Moreover, the computation of the likelihood of a network requires estimating probabilities of the observed data to realise during the evolution of the epidemiological process on this network. A stochastic simulation algorithm struggles to estimate rare events of observing the data (corresponding to the ground truth network) during the evolution with a significantly different network, and hence prevents optimisation of the likelihood. We replace the stochastic simulation with solving the chemical master equation for the probabilities of all network states. This equation also suffers from the curse of dimensionality due to the exponentially large number of network states. We overcome this by approximating the probability of states in the Tensor-Train decomposition. This enables fast and accurate computation of small probabilities and likelihoods. Numerical simulations demonstrate efficient black–box Bayesian inference of the network.
Keywords: epidemiological modelling, networks, tensor train, stochastic simulation algorithm, Markov chain Monte Carlo, Bayesian inference
2020 MSC: 15A69, 34A30, 37N25, 60J28, 62F15, 65F55, 90B15, 95C42
@@ 1 Introduction
The recent outbreak of COVID-19 and the public discussion that followed has led to better understanding of the central role that epidemiological models play in the decision making process and developing an informed response strategy. The quality of the mathematical models used in this process is crucial, not only because it allows to accurately predict the spread of a disease in population, but also in order to increase public trust in research and the decisions that follows from it. The limitations of the the over–simplified homogeneous models [48] in describing the human communities are well understood. To move beyond the Kermack–McKendrick compartmental SIR model [48], we should discard the assumption that population is ‘well–mixed’ and include the information on the structure of contact network, leading to study of epidemics on networks [47, 13, 75, 62, 49].
Predicting the evolution of epidemic on a given network is much more difficult than solving a homogeneous model. If the number of individuals assigned to each network node is small (which is the most interesting alternative to homogeneous models), the evolution becomes stochastic, and one needs to model a probability distribution function of states instead of states themselves. This probability distribution function satisfies the chemical master equation (CME) [71], which is an ordinary differential equation on the probability values. However, the total number of network states (and hence the size of the CME) grows exponentially in the number of network nodes [13, 75]. This makes the direct solution of the stochastic network models computationally intractable for large networks [49].
@@
If the contact network is not known, we can attempt to solve an inverse problem, i.e. to infer the network from observations of disease data over time. For network nodes, the number of possible networks grows exponentially in Hence, for large the problem complexity typically grows much faster than the information available, and network inference becomes a (very) underdetermined problem [16]. Network inference is therefore only solved directly for very small population sizes [59, 24]. Networks with mass-action kinetics can be inferred uniquely by observing transition rates at a simplex set of states [25], but the cardinality of this set is combinatorial in .
To address the problem for larger one can involve additional information about the network structure, such as degree distribution and sparsity [57, 35]. In a similar way, one can assume the underlying statistical distribution for the network and infer its parameters [33]. Another approach is to infer properties of the network rather than its exact structure [9], e.g. to find a class of network distributions, which the contact network most likely belongs to [17]. Related work include inferring the origin of epidemic given the contact network [53]. For a recent survey on network inference see [10].
In this paper we investigate inference of the contact network from states of the network observed over time. We use a Bayesian formulation but do not assume any prior knowledge on the structure of the contact network, and thus solve a black–box optimisation problem
i.e. look for the optimal network for which the actually observed data are most likely to appear. To summarise the above, the network inference problem is difficult due to the following reasons.
@@
Next, we apply two Markov Chain Monte Carlo (MCMC) algorithms for black–box discrete high–dimensional optimisation and analyse results. An overall workflow of this procedure is illustrated in Figure 1. By simulating three examples of networks, we show that by collecting sufficiently many data we can make the contrast high enough to infer the original ground truth network.
@@ 2.4 Tensor product algorithms for forward problem
for Let us now uncover the tensor product structure of the matrix of this CME, assuming that the probability of the network state appears in the probability distribution function (p.d.f.) vector in position
@@
we can write and Collecting these reaction rates in vectors of size , and using the big–endian ordering as explained above, we obtain tensor product decompositions
where appears in position appear elsewhere; and
@@
This so-called canonical polyadic (CP) [41, 36, 12] form represents the CME matrix as a sum of elementary tensors, each of which is a direct product of small matrices, acting on a single node of the network only. Hence, the storage for reduces from down to elements, where denotes the average degree of The curse of dimensionality for the matrix is therefore removed.
@@ 2.5 Ordering of network nodes for faster forward problem solving
These geometric properties of the network can be estimated from the eigenvector also known as Fiedler vector. In the pioneering paper [27] it was related to finding an optimal cut in the network It was generalised to reducing the envelope (or the bandwidth) of the adjacency matrix in [6], and later to finding orderings of variables for which TT decomposition (12), and related MPS and DMRG representations in quantum physics [66], have lower TT ranks [5]. The Fiedler vector can be computed by minimising the Rayleigh quotient also known as the Courant minimax principle, orthogonally to Following [6], we use the Fiedler vector to define a one–dimensional embedding of the graph to a linear chain. Let be the permutation vector of the set of nodes such that is ordered, i.e. or equivalently in the ascending order. Following [5], the same permutation of variables also reduces the TT ranks of the TT decomposition (12). In particular, it groups together variables corresponding to independent subgrids. Hence, we compute this permutation and adopt its order of variables before solving the forward problem (10) using tensor product algorithms.
@@ 2.6 Algorithms for Bayesian inverse problem
Since the full grid search is unfeasible for even moderate networks, we adopt the Metropolis–Hastings Markov Chain Monte Carlo (MCMC) method to approach the maximum of . The method is depicted in Algorithm 1. We need to choose a proposal distribution which is tractable for sampling a new state of the network , given the current network . In each iteration, given the current network , we sample a new proposal and accept or reject it probabilistically based on the Metropolis–Hastings ratio, forming a Markov Chain of network configurations After the Markov Chain is computed, we return the sample of this chain with the maximal likelihood.
This algorithm is known to converge to the true distribution (where is the normalising constant) under mild assumptions [64]. In practice, we use two proposal distributions.
@@
Choose one link in uniformly at random and toggle its state (on/off). Since there are possible links to toggle, independently of and , and hence cancels in the Metropolis-Hastings ratio. We will call Algorithm 1 with this proposal “MCMC”.
@@
Every iterations (i.e. when ), sample a random permutation vector of the set and in the next iterations toggle links in the order prescribed by . This is still a valid MCMC algorithm with a constant proposal distribution, but now with respect to , corresponding to a collection of networks in consecutive update steps, . However, as we will observe in the numerical experiments, this algorithm can often increase the likelihood faster in terms of the individual link changes, and hence the computing time. In each block of iterations, this algorithm proposes link changes without replacement [67]. For this reason, we will call this algorithm “MCMC noR” (for “no Replacement”).
Cancelling the constant proposal probability, and using log-likelihoods to avoid numerical over- and under-flow errors, we arrive at the following formula for the Metropolis–Hastings ratio:
@@ 3 Results
The numerical experiments were implemented in Matlab 2022b based on TT-Toolbox111https://github.com/oseledets/TT-Toolbox and tAMEn222https://github.com/dolgov/tamen packages, and run on one node of the HC44 series of the University of Bath “Nimbus” Microsoft Azure cluster, parallelising experiments with different datasets over cores of the Intel Xeon Platinum 8168 CPUs. Each of these parallel processes ran in a single–threaded mode.
The codes are publicly available from github.com/savostyanov/ttsir.
@@ 3.1 Linear chain
First, we checked the contrast of the log–likelihood at the ground truth network by computing averaged over the sampled datasets. We calculate this value for all that are nearest neighbours of i.e. differ by only one link. The results are shown on Fig. 4(b). We note that removal of an existing link from results in a contrast raising to for links attached to the sides of the chain. This is easy to understand, as removal of a link from creates a disconnected graph where two subraphs can not pass the infection on to each other, hence the epidemic dynamics on differs significantly from the one on Adding a new link to results in a milder contrast because the grid remains connected and the dynamics of the epidemic is less affected. This confirms that is at least a local optimum for and therefore can be inferred by Bayesian optimisation, assuming the optimisation algorithm manages to converge to it.
@@
We performed the black–box Bayesian optimisation using steps of the MCMC algorithms with and without replacement as explained in Sec. 2.6. The results are shown in Fig. 4(d) for the convergence of the likelihood towards the one of the ground truth network, and in Fig. 4(e) for the corresponding convergence of the network towards the ground truth network The latter is measured using the number of incorrectly inferred links,
related to the total number of possible links, With both algorithms we observe a steady convergence towards optimum with the ground truth grid correctly inferred in out of experiments and one link inferred incorrectly in out of experiments after likelihood evaluations. In total for this experiment, the network inference from each dataset with records took about seconds.
@@ 3.2 Austria road network
For this experiment we considered a more realistic example of a contact network, drawn from the road network in Austria, shown in Fig. 5(a). From preliminary experiments we noted that both MCMC algorithms for Bayesian optimisation struggle to converge to the optimum. To partly mitigate this, we increased the size of the dataset. We generated samples of synthetic data for a –SIS model with per contact transfer rate individual recovery rate and self–infection rate for the duration of time units, and again re-sampled the data to a uniform grid with the time step hence creating data records for each data sample.
@@
Using the tensor product approach with the same parameters as in Sec. 3.1 for the forward problem, we were able to resolve probabilities of up to which produced non–zero values for all likelihoods , enabling the optimisation for the inverse problem. For this example, one likelihood evaluation solving the forward problem with tAMEn took seconds for and seconds for The main reason for the larger times compared to the previous experiment is the larger data size compared to for the linear chain. However, a more complex structure of the contact network also contributed via higher TT ranks for and for The total time required to perform the Bayesian optimisation with steps of the MCMC algorithm took us about seconds, or just slightly over 2 days. For comparison, when SSA is used as the forward solver, one likelihood computation took seconds for and seconds for
From results shown in Fig. 5 we note that the convergence of both MCMC algorithms is stuck at a level where a small number of links of the network are inferred incorrectly, namely, approximately of links using the standard MCMC algorithm, and approximately of links using the MCMC algorithm without replacements. This highlights the potential of the no-replacement MCMC algorithm to tackle more complex problems, but also motivates the development of alternative approaches to black–box high–dimensional optimisation to improve their convergence and hence the quality of inference.
@@
The fact that the MCMC algorithms got stuck before the ground state network is inferred can be explained by presence of local maxima in which the MCMC methods get trapped. This can be considered as a consequence of a too high contrast, that sharpens the high–dimensional landscape and makes both the global and local maxima steeper. If we use less data for Bayesian inference, the contrast reduces, making it easier for MCMC to escape from local optima by switching from current network to less attractive proposal with probability as explained in Alg. 1. However, it also makes global maximum less emphasized and can lead to a situation where the optimal grid recovered by the Bayesian optimisation is not the same as the ground truth grid,

### v2

@@ Abstract
We consider a problem of inferring contact network from nodal states observed during an epidemiological process. In a black–box Bayesian optimisation framework this problem reduces to a discrete likelihood optimisation over the set of possible networks. The cardinality of this set grows combinatorially with the number of network nodes, which makes this optimisation computationally challenging. For each network, its likelihood is the probability for the observed data to appear during the evolution of the epidemiological process on this network. This probability can be very small, particularly if the network is significantly different from the ground truth network, from which the observed data actually appear. A commonly used stochastic simulation algorithm struggles to recover rare events and hence to estimate small probabilities and likelihoods. In this paper we replace the stochastic simulation with solving the chemical master equation for the probabilities of all network states. Since this equation also suffers from the curse of dimensionality, we apply tensor train approximations to overcome it and enable fast and accurate computations. Numerical simulations demonstrate efficient black–box Bayesian inference of the network.
keywords:
epidemiological modelling, networks, tensor train, stochastic simulation algorithm, Markov chain Monte Carlo, Bayesian inferencepacs:
[MSC Classification] 15A69, 34A30, 37N25, 60J28, 62F15, 65F55, 90B15, 95C42
@@ 1 Introduction
1.1 Background
The recent outbreak of COVID-19 and the public discussion that followed has led to better understanding of the central role that epidemiological models play in the decision making process and developing an informed response strategy. The quality of the mathematical models used in this process is crucial, not only because it allows to accurately predict the spread of a disease in population, but also in order to increase public trust in research and the decisions that are based on it. A large number of epidemic models used in education and research follow the Kermack–McKendrick compartmental SIR model [1]. An implicit assumption of this model is that the susceptible, infected and recovered groups are well–mixed in the sense that every person in any group has the same probability to come in contact with anyone from another group, i.e. the contact network is homogeneous. However, this assumption holds relatively poorly in communities, which poses a significant limitation to application of compartmental models. More accurate models should include the information on the structure of contact network, leading to study of epidemics on networks [2, 3, 4, 5, 6].
Predicting the evolution of epidemic on a given network is much more difficult than solving a homogeneous model. Consider the case when each network node represents a single individual, who can be susceptible or infected. The dynamics of the epidemic becomes stochastic and depends on the position of infected individuals in the network and the number of susceptible neighbours they are in contact with. Hence, we need to model a probability distribution function of network states instead of states themselves. This probability distribution function satisfies the chemical master equation (CME) [7], which is an ordinary differential equation on the probability values. However, the total number of network states (and hence the size of the CME) grows exponentially in the number of network nodes [3, 4]. This makes the direct solution of the stochastic network models computationally intractable for large networks [6].
@@ 1.1 Background
If the contact network is not known, we can attempt to solve an inverse problem, i.e. to infer the network from observations of disease data over time. For network nodes, the number of possible networks grows exponentially in Hence, for large the problem complexity typically grows much faster than the information available, and network inference becomes a (very) under-determined problem [19]. Network inference is therefore only solved directly for very small population sizes [20, 21], or equivalently by assuming that the population consists of a few densely connected groups and estimating couplings between them [22]. Networks with mass-action kinetics can be inferred uniquely by observing transition rates at a simplex set of states [23], but the cardinality of this set is combinatorial in . To address the problem for larger one can involve additional information about the network structure, such as degree distribution and sparsity [24, 25], community structure [26], and/or assumptions on the underlying statistical distribution for the network and infer its parameters [27]. In some cases more complex than pairwise interactions improve the network reconstruction [28]. Predictability of a stochastic process of observations and reconstructability of a network using their mutual information was considered in [29]. This information can be used to estimate the success of a network inference, before the full inference algorithm is applied. In contrast to stochastic inference, one can instead model a deterministic dynamics and minimise the observational error over the network parameters in the right-hand side of the dynamics [30]. Another approach is to infer properties of the network rather than its exact structure [31], e.g. to find a class of network distributions, which the contact network most likely belongs to [15]. Related work include inferring the origin of epidemic given the contact network [32]. For a recent survey on network inference see [33]. Finally, most close to our work is the maximum-likelihood estimation of the network link probabilities from binary time series [34]. However, the latter paper uses an expectation-maximisation algorithm assuming the Poisson distribution, while in this paper we rely on the full Bayesian formalism with likelihoods computed directly from the chemical master equation.
1.2 Our contribution
In this paper we investigate inference of the contact network from states of the network observed over time. We use a Bayesian formulation to find the network with the maximum a posteriori (MAP) estimate. However, we do not assume any prior knowledge on the structure of the contact network (uniform prior), and thus look for the maximum likelihood estimate (MLE),
which we solve as black–box optimisation problem. To summarise the above, the network inference problem is difficult due to the following reasons.
@@ 1.2 Our contribution
Next, we apply two Markov Chain Monte Carlo (MCMC) algorithms for black–box discrete high–dimensional optimisation and analyse results. An overall workflow of this procedure is illustrated in Fig. 1. By simulating three examples of networks, we show that by collecting sufficiently many data we can make the contrast high enough to infer the original ground truth network.
1.3 Notation
We use calligraphic font for contact networks blackboard font for sets sans-serif font for probability , expectation variance and likelihood Indices and scalar values are shown in usual maths italics. We use maths roman font for vectors, e.g. network states or unit vectors We use capitals for matrices, e.g. adjacency matrix of network When the considered vectors and matrices grow exponentially in number of people and hence suffer from the curse of dimensionality, we highlight it using bold font, e.g. for high–dimensional probability distribution function, and for the matrix of CME.
@@ 2.4 Tensor product algorithms for forward problem
for Let us now uncover the tensor product structure of the matrix of this CME. For this we assume that network states are ordered lexicographically, i.e. the state has index
which corresponds to how binary numbers are written in big–endian notation, e.g. see also Fig. 4. When the probability distribution function (p.d.f.) vector is composed, we place the probability in position , assuming 0-based indexing scheme, i.e. vector indices enumerated from up to If 1-based indexing scheme is used, we place in position
@@
we can write and Collecting these reaction rates in vectors of size , and using the big–endian lexicographic ordering as explained above, we obtain tensor product decomposition
where appears in position appear elsewhere. The tensor product structure is illustrated for in Fig. 4. For example, the full vector of recovery transitions corresponding to recovery of person is
We can see that this transition is possible from states with which are located in positions of the vector (assuming 0-based indexing), in agreement to eq. (8) and Fig. 4(a).
Similarly,
@@
where and This so-called canonical polyadic (CP) [61, 62, 63] form represents the CME matrix as a sum of elementary tensors, each of which is a direct product of small matrices, acting on a single node of the network only. Hence, the storage for reduces from down to elements, where denotes the average degree of The curse of dimensionality for the matrix is therefore removed.
@@ 2.5 Ordering of network nodes for faster forward problem solving
These geometric properties of the network can be estimated from the eigenvector also known as Fiedler vector. In the pioneering paper [71] it was related to finding an optimal cut in the network It was generalised to reducing the envelope (or the bandwidth) of the adjacency matrix in [72], and later to finding orderings of variables for which TT decomposition (13), and related MPS and DMRG representations in quantum physics [73], have lower TT ranks [69]. The Fiedler vector can be computed by minimising the Rayleigh quotient also known as the Courant minimax principle, orthogonally to Following [72], we use the Fiedler vector to define a one–dimensional embedding of the graph to a linear chain. Let be the permutation vector of the set of nodes such that is ordered, i.e. or equivalently in the ascending order. Following [69], the same permutation of variables also reduces the TT ranks of the TT decomposition (13). In particular, it groups together variables corresponding to independent subnetworks. Hence, we compute this permutation and adopt its order of variables before solving the forward problem (11) using tensor product algorithms.
@@ 2.6 Algorithms for Bayesian inverse problem
Since the full grid search is unfeasible for even moderate networks, we adopt the Metropolis–Hastings Markov Chain Monte Carlo (MCMC) method to approach the maximum of . The method is depicted in Algorithm 1. We need to choose a proposal distribution which is tractable for sampling a new state of the network , given the current network . In each iteration, given the current network , we sample a new proposal and accept or reject it with probability based on the Metropolis–Hastings ratio, forming a Markov Chain of network configurations After the Markov Chain is computed, we return the sample of this chain with the maximal likelihood.
This algorithm is known to converge to the true distribution (where is the normalising constant) under mild assumptions [74]. We implement two proposal distributions.
@@
Choose one link in uniformly at random and toggle its state (on/off). Since there are possible links to toggle, independently of and , and hence cancels in the Metropolis-Hastings ratio. We will call Algorithm 1 with this proposal “MCMC-R”, since it samples the links with Replacement.
@@
Every iterations (i.e. when ), sample a random permutation vector of the set and in the next iterations toggle links in the order prescribed by . This is still a valid MCMC algorithm with a constant proposal distribution, but now with respect to , corresponding to a collection of networks in consecutive update steps, . In our numerical experiments reported in Section 3 this algorithm sometimes increased the likelihood faster in terms of the individual link changes (and hence computing time), and resulted in a better grid reconstruction. In each block of iterations, this algorithm proposes link changes without replacement [75]. For this reason, we will call this algorithm “MCMC-noR” (for “no Replacement”).
Cancelling the constant proposal probability, and using log-likelihoods to avoid numerical over- and under-flow errors, we can rewrite the Metropolis–Hastings ratio as We can further modify this formula to gain more control on performance of MCMC. Specifically, we introduce the temperature, or tempering parameter as follows:
We note that this formula resembles the Boltzmann distribution, also know as Gibbs distribution, which is used in statistical and theoretical physics and describes probability to observe particles in the ground and exited energy states. Similarly, the parameter controls how often MCMC accepts a network with worse likelihood that the current one, which in order affects the convergence of the algorithm and its ability to get out of local optima. If a new grid has better likelihood than the current grid it is always accepted. Otherwise, it is accepted only with probability for . For this probability becomes which increases the probability that the new grid is accepted, encouraging MCMC to escape a local maximum.
@@ 3 Results
The numerical experiments were implemented in Matlab 2022b based on TT-Toolbox222https://github.com/oseledets/TT-Toolbox and tAMEn333https://github.com/dolgov/tamen packages, and run on one node of the HC44 series of the University of Bath “Nimbus” Microsoft Azure cluster. Experiments with different datasets were run in parallel over cores of the Intel Xeon Platinum 8168 CPUs. Each of these parallel processes ran in a single–threaded mode. The codes are made public and freely available from github.com/savostyanov/ttsir.
@@ 3.1 Linear chain
First, we checked the contrast of the log–likelihood at the ground truth network by computing for all that are nearest neighbours of i.e. differ by only one link. The results are averaged over the sampled datasets and shown in Fig. 5(b). We note that removal of an existing link from results in contrast raising to for links attached to the sides of the chain. This is easy to understand, as removal of a link from creates a disconnected network where two parts can not pass the infection on to each other, hence the epidemic dynamics on differs significantly from the one on Adding a new link to results in a milder contrast because the grid remains connected and the dynamics of the epidemic is less affected. This confirms that is at least a local optimum for and therefore can be inferred by Bayesian optimisation, assuming the optimisation algorithm manages to converge to it.
@@
We performed the black–box Bayesian optimisation using steps of the MCMC algorithms with and without replacement as explained in Sec. 2.6. We observed similar performance of both algorithms, hence only the results for MCMC-noR are shown in Fig. 5(e) for the convergence of the likelihood towards the one of the ground truth network, and in Fig. 5(f) for the corresponding convergence of the network towards the ground truth network The latter is measured using the number of incorrectly inferred links,
related to the total number of possible links, For both MCMC-R and MCMC-noR without tempering (with ), we observe a steady convergence towards optimum with the ground truth grid correctly inferred in out of experiments and one link inferred incorrectly in out of experiments after likelihood evaluations. To improve this result, we used tempering with temperature and observed a slightly slower convergence of MCMC-noR, which then achieved the exact recovery for all data samples. We also observed that increasing the temperature further to results in a much slower convergence and poor recovery, indicating that this parameter needs to be carefully adjusted. In total for this experiment, the network inference from each dataset with records took about seconds.
@@ 3.2 Austria road network
For this experiment we considered a more realistic example of a contact network, drawn from the road network in Austria, shown in Fig. 6(a). As previously, we generated samples of synthetic data for a –SIS model with per contact transfer rate individual recovery rate and self–infection rate However, from preliminary experiments we noted that both MCMC algorithms for Bayesian optimisation struggle to converge to the optimum. To partly mitigate this, we increased the size of each dataset to data records, created by observing the state for the duration of time units at uniform time grid with the step
@@
Using the tensor product approach with the same parameters as in Sec. 3.1 for the forward problem, we were able to resolve probabilities of up to which produced non–zero values for all likelihoods , enabling the optimisation for the inverse problem. For this example, one likelihood evaluation solving the forward problem with tAMEn took seconds for and seconds for The main reason for the larger times compared to the previous experiment is the larger data size compared to for the linear chain. However, a more complex structure of the contact network also contributed via higher TT ranks for and for The total time required to perform the Bayesian optimisation with steps of the MCMC algorithm took us about hours. For comparison, when SSA is used as the forward solver, one likelihood computation took seconds for and seconds for
From results shown in Fig. 6 we note that without tempering () the convergence of both MCMC algorithms is stuck in a local maximum where approximately of links are inferred incorrectly. Using tempering with we observed almost the same convergence at initial stage of optimisation, which then resulted in a faster convergence towards a better inference, with out of data samples allowed for the exact recovery, and the remaining out of had only one incorrect link out of
@@
The difference in performance compared to the linear chain example can be considered as a consequence of a high contrast, that sharpens the high–dimensional landscape and makes both the global and local maxima steeper. If we use less data for Bayesian inference, the contrast reduces, making it easier for MCMC to escape from local optima by switching from current network to less attractive proposal with probability as explained in Alg. 1. However, it also makes global maximum less emphasised and can lead to a situation where the optimal grid recovered by the Bayesian optimisation is not the same as the ground truth grid,
3.3 Florentine families
We consider a slightly larger network representing marriage alliances and business relationships between Florentine families in XV century.444The network is taken from networks.skewed.de/net/florentine_families. The network is given as an undirected weighted graph with nodes, one of which is isolated from the rest, as shown in Figure 7(a). The weights of the links represent what kind of relationship the families have. For the purpose of this experiment we ignore the disconnected node and disregard the difference in connections. Hence we consider an undirected and unweighted network of nodes. Still, a more densely connected network leaves states more frequently in the infected state with used in the previous examples. To obtain a more meaningful data (and more accurate inference), we simulate the observation data using , , and . We observe states at uniformly distributed time points over the duration of time units, sampled with the time step resulting in data records.
From the contrasts shown in Figure 7(b) we see that removal of the link between nodes and results in disconnected network, which results in a significant contrast, Notably, removal of the link does not fully disconnect the network, but considerably reduces the chance for the disease to reach from the node to the node hence the observed large contrast Similarly, removal of the link makes it harder for the disease to reach the node which also results in a high contrast The remaining links seem to be considerably less important, and their removal results in a lower contrast On average, adding extra links also result in a lower contrast, with a notable exception of the first node, which is the origin of the epidemic. The lower contrast around the ground truth network may cause extra challenge for the exact recovery of this network, particularly if the likelihoods are computed inaccurately.
We run the MCMC-noR algorithm without tempering (), and with tempering (), but observe better results of the former. The convergence of the log-likelihood is shown in Figure 7(c), and the convergence of the inferred network is shown in Figure 7(d). We observe accurate recovery of the network, specifically, among the data samples that we tried, resulted in exact recovery. In the remaining data samples, we observed only to incorrectly recovered links. In our experiments, the MCMC-noR algorithm has reached the final network configuration after samples on average, which took days of CPU time.